# An Image is Worth More Than 16$\times$16 Patches: Exploring Transformers on Individual Pixels

### v1

@@ Discussion.
As another way to remove locality, pixel permutation is highly destructive. On the other hand, PiT shows successful elimination of locality is possible by treating individual pixels as tokens. We hypothesize this is because permuting pixels not only damages the locality bias, but also hurts the other inductive bias – translation equivariance. In PiT, although locality is removed altogether, the Transformer weights are still shared to preserve translation equivariance; but with shuffling, this inductive bias is also largely removed. The difference suggests that translation equivariance remains important and should not be disregarded, especially after locality is already compromised.
$T$, $\delta{=}\inf$ Acc@1 $\Delta$Acc
@@
Figure 5: Results of pixel permutation for ViT-B on ImageNet. We vary the number of pixel swaps $T$ (left) and maximum distance of pixel swaps $\delta$ (right). Pixel permutations can drop accuracy by 25.2% (when 25K pairs are swapped), compared to the relatively minor drop (1.6%) when the position embedding is completely removed. These results suggest pixel permutation imposes a much more significant impact on performance, compared to swapping position embeddings.
@@ 7 Conclusion and Limitations
Through our explorations, we have demonstrated that Transformers can directly work with individual pixels as tokens. This is surprising, as it allows for a clean, potentially scalable architecture without locality – an inductive bias that was presumably fundamental for vision models. Given the spirit of deep learning that aims to replace manually inducted priors with data-driven, learnable alternatives, we believe our finding is of great value to the community, especially when designing the next-generation of models for the domain of 2D images and beyond.
@@
However, the practicality and coverage of our current demonstrations remains limited. Given the quadratic computation complexity, PiT is more of a method for investigation, and less for applications. And even with the additional tasks in Appendix 0.C, the study is still not comprehensive. Nonetheless, we believe this work has sent out a clear, unfiltered message that locality is not fundamental, and patchification is simply a useful heuristic that trades-off efficiency vs. accuracy.
Acknowledgments.
We thank Kaiming He, Mike Rabbat and Sho Yaida for helpful discussions. We thank Yann LeCun for feedback on positioning.
Appendix 0.A Visualizations of PiT
@@
To check what PiT has learned, we experimented different ways for visualizations. Unless otherwise specified, we use PiT-B and ViT-B models trained with supervised learning on ImageNet classification, and compare them side by side.
@@
In Fig. 6, we present the mean attention distances for PiT and ViT across three categories: late layers (last 4), middle layers (middle 4), and early layers (first 4). Following [22], this metric is computed by aggregating the distances between a query token and all the key tokens in the image space, weighted by their corresponding attention weights. It can be interpreted as the size of the ‘receptive field’ for Transformers. The distance is normalized by the image size, and sorted based on the distance value for different attention heads from left to right.
As shown in Fig. 6(a) and Fig. 6(b), both models exhibit similar patterns in the late layers, with the metric increasing from the 8th to the 11th layer. In the middle layers, while ViT displays a mixed trend among layers (see Fig. 6(d)), PiT clearly extract patterns from larger areas in the relatively later layers (see Fig. 6(c)). Most notably, PiT focuses more on local patterns by paying more attention to small groups of pixels in the early layers, as illustrated in Fig. 6(e) and Fig. 6(f).
@@
Fig. 7 shows the mean attention offsets between PiT and ViT as introduced in [63]. This metric is calculated by determining the center of the attention map generated by a query and measuring the spatial distance (or offset) from the query’s location to this center. Thus, the attention offset refers to the degree of spatial deviation of the ‘receptive field’ – the area of the input that the model focuses on – from the query’s original position. Note that different from ConvNets, Self-Attention is a global operation, not a local operation that is always centered on the current pixel (offset always being zero).
Interestingly, Fig. 7(e) suggests that PiT captures long-range relationships in the first layer. Specifically, the attention maps generated by PiT focus on regions far away from the query token – although according to the previous metric (mean attention distance), the overall ‘size’ of the attention can be small and focused in this layer.
@@
In Fig. 8, we observe another interesting behavior of PiT. Here, we use the central pixel in the image space as the query and visualize its attention maps in the early layers. We find that the attention maps in the early layers can already capture the foreground of objects. Figure-ground segmentation [3] can be effectively performed with low-level signals (e.g., RGB values) and therefore approaches with a few layers. And this separation prepares the model to potentially capture higher-order relationships in later layers.
(a) PiT in late layers.
(b) ViT in late layers.
(c) PiT in middle layers.
(d) ViT in middle layers.
(e) PiT in early layers.
(f) ViT in early layers.
Figure 6: Mean attention distances in late, middle, and early layers between PiT and ViT. This metric can be interpreted as the receptive field size for Transformers. The distance is normalized by the image size, and sorted based on the distance value for different attention heads from left to right.
(a) PiT in late layers.
(b) ViT in late layers.
(c) PiT in middle layers.
(d) ViT in middle layers.
(e) PiT in early layers.
(f) ViT in early layers.
Figure 7: Mean attention offsets in late, middle, and early layers between PiT and ViT. This metric measures the deviation of the attention map from the current token location. The offset is normalized with the image size, and sorted based on the distance value for different attention heads from left to right.
Figure 8: Figure-ground segmentation in early layers of PiT. In each row, we show the original image and the attention maps of two selected early layers (from first to fourth). We use the central pixel in the image space as the query and visualize its attention maps. Structures that can capture the foreground of objects have already emerged in these layers, which prepares for the later layers to learn higher-order relationships.
@@
Table 6: Extended results for image generation. We continue the training from 400 epochs (main paper) to 1400 epochs, and find the gap between DiT and PiT becomes larger, especially on FID.
Appendix 0.B Extended Results on Image Generation
@@
In the main paper (Sec. 5.3), both image generation models, DiT-L/2 and PiT-L, are trained for 400 epochs. To see the trend for longer training, we followed [51] and simply continued training them till 1400 epochs while keeping the learning rate constant.
@@
The results are summarized in Tab. 6. Interestingly, longer-training also benefits PiT more than DiT. Note that FID shall be compared in a relative sense – a 0.2 gap around 2 is bigger than 0.2 around 4.
Appendix 0.C Generalization to Other Tasks
To further examine the generalization of our observation, we tried PiT on two more tasks: (i) fine-grained classification on Oxford-102-Flower [49], which requires nuanced understanding; and (ii) depth estimation on NYU-v2 [59]. Given the computation budget, we resize images to either 32$\times$32 (former) or 48$\times$64 (latter), and follow standard protocols to train and evaluate models. The results again shows PiT holds more effectiveness over ViT in quality:
fine-grained classification depth estimation
Acc@1 ($\uparrow$) Acc@5 ($\uparrow$) RMSE ($\downarrow$)
ViT-S/2 45.8 68.3 0.80
PiT-S 46.3 68.9 0.72
Appendix 0.D Texture vs. Shape Bias Analysis
@@ Appendix 0.C Generalization to Other Tasks
As a final interesting observation, we used an external benchmark [^4] https://github.com/rgeirhos/texture-vs-shape which checks if an ImageNet classifier’s decision is based on texture or shape. ConvNets are heavily biased toward texture ($\sim$20 in shape bias). Interestingly, we find PiT relies more on shape than ViT (57.2 vs. 56.7), suggesting that even when images are broken down into sets of pixels, Transformers can still sift through potentially abundant texture patterns to identify and rely on the sparse shape signals for tasks like object recognition.
Appendix 0.E Additional Notes for Training PiT
@@
While for some cases (e.g., DiT [51]), the training recipe can be directly transferred to PiT; for some other cases, we do want to note more potential challenges during training. Below we want to especially highlight the effect of reduced learning rates when training a supervised ViT from scratch.
We take CIFAR-100 as a representative example. As shown in Fig. 9, we find for PiT, the training becomes unstable if we maintain the same learning rate from ViT. It is especially vulnerable toward the end of the schedule. When the initial learning rate is reduced from $2e^{-3}$ to $1e^{-3}$, the training is more stable and leads to better accuracy. Similar observations are also made on ImageNet.
@@
Figure 9: Training curves of different learning rates (PiT-T, AdamW, batch 1024, CIFAR-100). We find with a reduced learning rate compared to ViT training, the curve is more stable, and ultimately leads to better accuracy.

### v2

@@ Discussion.
As another way to remove locality, pixel permutation is highly destructive. On the other hand, we show successful elimination of locality is possible by treating individual pixels as tokens. We hypothesize this is because permuting pixels not only damages the locality bias, but also hurts the other inductive bias – location equivariance. Although locality is removed, the Transformer weights are still shared across pixels to preserve location equivariance; but with shuffling, this inductive bias is also largely removed. The difference suggests that proper weight-sharing remains important and should not be disregarded, especially after locality is already compromised.
[…]
@@
Figure 5: Results of pixel permutation for ViT-B on ImageNet. We vary the number of pixel swaps $T$ (left) and additionally varying the maximum distance of pixel swaps $\delta$ (right). Pixel permutations can drop accuracy by 25.2% (when 25K pairs are swapped), compared to the relatively minor drop (1.6%) when the position embedding is completely removed. And when farther-away pixels are allowed for swapping, more damage is caused. These results suggest pixel permutation imposes a much more significant impact on performance, compared to swapping position embeddings.
6 Related Work
Locality for images.
To of our knowledge, most modern vision architectures (He et al., 2016; Parmar et al., 2018), including those aimed at simplifications of inductive biases (Dosovitskiy et al., 2021; Tolstikhin et al., 2021), still maintain locality in the design. Manual features before deep learning are also locally biased. For example, SIFT (Lowe, 2004) uses a local descriptor to represent each point of interest; HOG (Dalal & Triggs, 2005) locally normalizes the gradient strengths to account for changes in illumination and contrast. Interestingly, with these features, bag-of-words models (Csurka et al., 2004; Lazebnik et al., 2006) were popular – analogous to the set-of-pixels explored in our work.
Locality beyond images.
The inductive bias of locality is widely accepted beyond modeling images. For text, a natural language sequence is often pre-processed with ‘tokenizers’ (Sennrich et al., 2015; Kudo & Richardson, 2018), which aggregate the dataset statistics for grouping frequently-occurring adjacent characters into sub-words. Before Transformers, recurrent neural networks (Mikolov et al., 2010; Hochreiter & Schmidhuber, 1997) were the default architecture for such data, which exploit temporal connectivity to process sequences step-by-step. For even less structured data (e.g. point clouds (Chang et al., 2015; Dai et al., 2017)), modern networks (Qi et al., 2017; Zhao et al., 2021) will resort to various sampling and pooling strategies to increase their sensitivity to the local geometric layout. In graph neural networks (Scarselli et al., 2009), nodes with edges are often viewed as locally connected, and information is propagated through these connections to farther-away nodes. Such designs make them particularly useful for analyzing social networks, molecular structures, etc. Despite its higher computational cost, Transformers with global Self-Attention is increasingly favored for real-world problems due to its powerful pattern-learning capabilities.
Other notable efforts.
  • To remove locality for ConvNets, Brendel & Bethge (2019) replaced all spatial convolutional filters in a ResNet (He et al., 2016) with $\times1$ filters. It has values in interpretability, but without inter-pixel communications the resulting network is substantially worse in performance. Our work instead uses Transformers, which are inherently built on set operations, with Self-Attention handling all-to-all communications; understandably, we obtain better results.
  • Before ViT gained popularity, iGPT (Chen et al., 2020a) was proposed to directly pre-train Transformers on pixels following their success on text (Radford et al., 2018; Devlin et al., 2019). In retrospect, iGPT is a locality-free model for self-supervised next- (or masked-) pixel prediction. But despite extensive scaling efforts, its performance still falls short compared to simple contrastive pre-training (Chen et al., 2020b) on ImageNet. Later, ViT (Dosovitskiy et al., 2014) re-introduced locality (e.g., via patchification) into the architecture, achieving impressive results on many benchmarks including ImageNet. This shift cemented $\times16$ patches as the standard for vision tasks. However, it remained unclear whether ViT’s performance gains were primarily due to higher resolution or patch-based locality. Through systematic analyses, our work puts a conclusive remark, pointing to resolution as the enabler for ViT, not locality.
  • Perceiver (Jaegle et al., 2021b; a) is another series of architectures that operate directly on pixels for images. Aimed at being modality-agnostic, Perceiver designs latent Transformers with cross-attention modules to tackle the efficiency issue when the input is high-dimensional. However, this design is not as widely adopted as plain Transformers, which have consistently demonstrated scalability across multiple domains (Brown et al., 2020; Dosovitskiy et al., 2021). We show Transformers can indeed work directly with pixels, and given the rapid development of Self-Attention implementations to handle massive sequence length (up to a million(Dao et al., 2022; Liu et al., 2023), efficiency may not be a critical bottleneck even counting all the pixels.
  • Transformer has also been shown to work well for distribution modeling of generalized signals, including pixels in Tulsiani & Gupta (2021).
  • Our work can also be viewed as extending sequence length scaling to the extreme (individual pixels for images). Longer sequence (or higher resolution) is generally beneficial, as evidenced in Chen et al. (2021b); Hu et al. (2022); Peebles & Xie (2023). However, all of them stopped short of reaching the extreme case that completely gets rid of locality.
  • Besides individual pixels, there has been sustained interest in learning flexible and data-driven patches for vision models via grouping (Zhang & Maire, 2020; Ke et al., 2022; Ma et al., 2023; Deng et al., 2024). While conceptually appealing – e.g. can reconcile the long-sequence issue faced with pixels, they are yet to show salient practical promise in this context. Nonetheless, they are related work and we point to this direction for interested readers.
@@ 7 Conclusion and Limitations
Through multiple case studies, we have demonstrated that Transformers can directly work with individual pixels. This is surprising, as it allows for a clean, potentially scalable architecture without locality – an inductive bias presumably fundamental for vision models. Given the spirit of deep learning that aims to replace manual priors with data-driven, learnable alternatives, we believe our finding is along the right direction and valuable to the community.
@@
However, the practicality and coverage of our current demonstrations remain limited. Given the quadratic computation complexity, treating pixel-based Transformer is more of an approach for investigation, and less for applications. And even with several tasks, the study is still not comprehensive. Nonetheless, we believe this work has sent out a clear, unfiltered message that locality is not fundamental, and patchification is simply a useful heuristic that trades-off efficiency vs. accuracy.
Appendix A Implementation Details
Below, we list the implementation details for the three case studies conducted in Sec. 4.
Case study #1.
For CIFAR-100, due to the lack of optimal settings even for ViT, we search for the recipe and report results using model sizes T(iny) and S(mall). We use the augmentations from the released demo of He et al. (2020), as we found more advanced augmentations (e.g., AutoAug (Cubuk et al., 2018)) not helpful in this case. All models are trained using AdamW (Loshchilov & Hutter, 2019) with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.95$. We use a batch size of 1024, weight decay of 0.3, drop path (Huang et al., 2016) of 0.1, and initial learning rate of 0.004 which we found to be the best for all the models. We use a linear learning rate warm-up of 20 epochs and cosine learning rate decay to a minimum of 1e-6. Our training lasts for 2400 epochs, compensating for the small size of the dataset.
On ImageNet, we closely follow the scratch training recipe from Touvron et al. (2021); He et al. (2022) for ViT. Due to our computation budget, images are crop-and-resized to $28{\times}28$ as the low-resolution inputs by default. The training batch size is 4096, initial learning rate is $1.6{\times}10^{-3}$, weight decay is 0.3, drop path is 0.1, and training length is 300 epochs. MixUp (Zhang et al., 2018) (0.8), CutMix (Yun et al., 2019) (1.0), RandAug (Cubuk et al., 2020) (9, 0.5), and exponential moving average (0.9999) are used.
[…]
(a) ViT/1 in late layers.
(b) ViT/2 in late layers.
(c) ViT/1 in middle layers.
(d) ViT/2 in middle layers.
(e) ViT/1 in early layers.
(f) ViT/2 in early layers.
Figure 6: Mean attention distances in late, middle, and early layers between ViT/1 and ViT/2. This metric can be interpreted as the receptive field size for Transformers. The distance is normalized by the image size, and sorted based on the distance value for different attention heads from left to right.
Case study #2.
We follow standard MAE and use a mask ratio of 75% and select tokens randomly. Given the remaining 25% of visible tokens, the model needs to reconstruct masked regions using pixel regression. Since there is no known default setting for MAE on CIFAR-100 (even for ViT), we search for recipes and report results using models of Tiny and Small sizes. The same augmentations as in He et al. (2022) are applied to the images during the pre-training for simplicity. All models are pre-trained using AdamW with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.95$. We follow all of the hyper-parameters in He et al. (2022) for the pre-training of 1600 epochs except for the initial learning rate of 0.004 and a learning rate decay of 0.85 (Clark et al., 2020). Thanks to MAE pre-training, we can fine-tune our model with a higher learning rate of 0.024. We also set weight decay to 0.02, layer-wise rate decay to 0.65, and drop path to 0.3, $\beta_{2}$ to 0.999, and fine-tune for 800 epochs. Other hyper-parameters closely follow the scratch training recipe for supervised learning in case study #1.
Case study #3.
We followed the settings for DiT training, with a larger batch size (2048) to make the training faster (the original recipe uses a batch size of 256). To make the training stable, we perform linear learning rate warm up (Goyal et al., 2017) for 100 epochs and then keep it constant for a total of 400 epochs. We use a maximum learning rate of 8e-4, with no weight decay applied. For generation, a sampling step of 250 is used.
Appendix B Visualizations of Transformer on Pixels
@@
To check what pixel-based Transformer has learned, we experimented different ways for visualizations. Unless otherwise specified, we use ViT-B and the locality-free variant (i.e., ViT-B/1) trained with supervised learning on ImageNet classification, and compare them side by side.
@@
In Fig. 6, we present the mean attention distances for ViT/1 and ViT across three categories: late layers (last 4), middle layers (middle 4), and early layers (first 4). Following Dosovitskiy et al. (2021), this metric is computed by aggregating the distances between a query token and all the key tokens in the image space, weighted by their corresponding attention weights. It can be interpreted as the size of the ‘receptive field’ for Transformers. The distance is normalized by the image size, and sorted based on the distance value for different attention heads from left to right.
As shown in Fig. 6(a) and Fig. 6(b), both models exhibit similar patterns in the late layers, with the metric increasing from the 8th to the 11th layer. In the middle layers, while ViT displays a mixed trend among layers (see Fig. 6(d)), ViT/1 clearly extract patterns from larger areas in the relatively later layers (see Fig. 6(c)). Most notably, ViT/1 focuses more on local patterns by paying more attention to small groups of pixels in the early layers, as illustrated in Fig. 6(e) and Fig. 6(f).
@@
Fig. 7 shows the mean attention offsets between ViT/1 and ViT as introduced in Walmer et al. (2023). This metric is calculated by determining the center of the attention map generated by a query and measuring the spatial distance (or offset) from the query’s location to this center. Thus, the attention offset refers to the degree of spatial deviation of the ‘receptive field’ – the area of the input that the model focuses on – from the query’s original position. Note that different from ConvNets, Self-Attention is a global operation, not a local operation that is always centered on the current pixel (offset always being zero).
Interestingly, Fig. 7(e) suggests that ViT/1 captures long-range relationships in the first layer. Specifically, the attention maps generated by ViT/1 focus on regions far away from the query token – although according to the previous metric (mean attention distance), the overall ‘size’ of the attention can be small and focused in this layer.
@@
In Fig. 8, we observe another interesting behavior of ViT/1. Here, we use the central pixel in the image space as the query and visualize its attention maps in the early layers. We find that the attention maps in the early layers can already capture the foreground of objects. Figure-ground segmentation (Boykov et al., 2001) can be effectively performed with low-level signals (e.g., RGB values) and therefore approaches with a few layers. And this separation prepares the model to potentially capture higher-order relationships in later layers.
(a) ViT/1 in late layers.
(b) ViT/2 in late layers.
(c) ViT/1 in middle layers.
(d) ViT/2 in middle layers.
(e) ViT/1 in early layers.
(f) ViT/2 in early layers.
Figure 7: Mean attention offsets in late, middle, and early layers between ViT/1 and ViT/2. This metric measures the deviation of the attention map from the current token location. The offset is normalized with the image size, and sorted based on the distance value for different attention heads from left to right.
Figure 8: Figure-ground segmentation in early layers of ViT/1. In each row, we show the original image and the attention maps of two selected early layers (from first to fourth). We use the central pixel in the image space as the query and visualize its attention maps. Structures that can capture the foreground of objects have already emerged in these layers, which prepares for the later layers to learn higher-order relationships.
@@
Table 6: Extended results for image generation. We continue the training from 400 epochs (main paper) to 1400 epochs, and find the gap between DiT/2 and DiT/1 becomes larger, especially on FID.
Appendix C Extended Results on Image Generation
@@
In the main paper (Sec. 4.3), both image generation models, DiT-L/2 and DiT-L/1, are trained for 400 epochs. To see the trend for longer training, we followed Peebles & Xie (2023) and simply continued training them till 1400 epochs while keeping the learning rate constant.
@@
The results are summarized in Tab. 6. Interestingly, longer-training also benefits DiT/1 more than DiT. Note that FID shall be compared in a relative sense – a 0.2 gap around 2 is bigger than 0.2 around 4.
Appendix D Texture vs. Shape Bias Analysis
@@
As a final interesting observation, we used an external benchmark [^3] https://github.com/rgeirhos/texture-vs-shape which checks if an ImageNet classifier’s decision is based on texture or shape. ConvNets are heavily biased toward texture ($\sim$20 in shape bias). Interestingly, we find ViT/1 relies more on shape than ViT (57.2 vs. 56.7), suggesting that even when images are treated as sets of pixels, Transformers can still sift through potentially abundant texture patterns to identify and rely on the sparse shape signals for tasks like object recognition.
Appendix E Additional Notes for Training Transformer on Pixels
@@
While for some cases (e.g., DiT (Peebles & Xie, 2023)), the training recipe can be directly transferred to pixel-based Transformer; for some other cases, we do want to note more potential challenges during training. Below we want to especially highlight the effect of reduced learning rates when training a supervised ViT from scratch.
@@
Figure 9: Training curves of different learning rates (ViT-T/1, batch size 1024, CIFAR-100). With a reduced learning rate from ViT/2, the curve is more stable, and ultimately leads to better accuracy.
We take CIFAR-100 as an example. As shown in Fig. 9, we find for pixel-based Transformer, the training becomes unstable if we maintain the same learning rate as vanilla ViTs. It is especially vulnerable toward the end. When the initial learning rate is reduced from $2e^{-3}$ to $1e^{-3}$, the training is more stable and leads to better accuracy. Similar observations are also made on ImageNet.
dictionary Acc@1 Acc@5
ViT-B/2 75.7 92.3
ViT-B/1 76.1 92.6
76.6 92.8
Table 7: Compared to the default ViT that uses parametric projection layers, we find a simple dictionary-based ViT that leverages the reduced vocabulary size to learn non-parametric embeddings for pixel intensity values works better for ImageNet supervised classification.
Appendix F Exploring Dictionary-Based ViT
As discussed in Sec. 3, one advantage for treating individual pixels as tokens for ViT is the reduction in vocabulary size. Specifically, the size of the vocabulary (i.e., the unique values the model must handle) becomes much more manageable.
To exploit this fact, we experimented with a dictionary-based approach for pixel tokens in ViT (other settings exactly follow default ViT for supervised learning). Specifically, we employed a non-parametric, 256-dimensional embedding for each of the 256 possible values a pixel can take, and different color channels have different embeddings. For a pixel with three (RGB) values, we mapped each value to its corresponding vector, and concatenated the three vectors. An optional linear projection was applied to ensure compatibility with the hidden dimension $d$ of ViT. Note that this approach has even less inductive bias because we no longer assume the relative order for pixel intensity values: the embedded value of 127 is not necessarily in-between 255 and 0, whereas it must be in-between if modeled directly with a parametric projection.
Tab. 7 summarizes the results. Interestingly, we find this ViT variant works better than our default ViT-B/1 on ImageNet. The improvement in performance further endorses our finding: 1) reducing inductive bias can be beneficial; and 2) locality-free ViTs can unlock the potential of dictionary-based token representations for next-generation vision architectures.
downsizing Acc@1 Acc@5
ViT-B/2 bicubic 75.7 92.3
ViT-B/1 76.1 92.6
ViT-B/2 single-pixel 70.5 89.0
ViT-B/1 75.6 92.3
Table 8: Influence of downsizing algorithms to our finding. Here we specifically experimented to downsize the original image by simply copying the value of its corresponding pixel from the original image (‘single-pixel’), as opposed to ‘bicubic interpolation’ which leverages nearby pixels/locality. While single-pixel downsizing are generally worse, locality-free models are much more robust to such a change for their representations being locality-agnostic.
Appendix G Influence of Downsizing Algorithms
By default, we use bicubic interpolation to perform downsizing. But since such algorithms also leverage nearby pixel values to estimate the value of the current pixel, there can be concerns that our pixel-based Transformers still rely on subtle locality biases.
To investigate the influence of downsizing algorithms, we have conducted additional experiments comparing bicubic interpolation with a simpler approach where each pixel in the downsized image directly copies the value of its corresponding pixel from the original image. In this way, it circumvents the locality bias introduced by interpolation. We call this downsizing algorithm ‘single-pixel’.
The results for ImageNet classification, ViT-B are summarized in Tab. 8. Two observations are made. First, as expected, single-pixel downsizing results in lower performance compared to bicubic interpolation. But second and more interestingly, patch-based ViT-B/2 shows a more significant performance drop than the locality-free ViT-B/1, and the gap becomes larger. This suggests that locality-based models are more sensitive to this change in downsizing algorithm. In contrast, locality-free models are more robust, likely due to their ability to learn locality-agnostic representations and do not rely on the inductive bias of locality for classification.
Appendix H Exploration of tokens via advanced grouping
Before our journey with pixels as tokens, we were interested in tokens via advanced grouping (i.e., SLIC (Achanta et al., 2010) and FH segmentation (Felzenszwalb & Huttenlocher, 2004)) due to its ability to quickly produce high-semantic and irregular shape tokens.
To compare these advanced tokenizations, we setup an experiment with the scratch training recipe from Touvron et al. (2021); He et al. (2022) with patch size being $16\times 16$ and sequence length of 196. For FH segmentation, we use the implementation from scikit. The optimal FH scale (which indirectly controls the resulting sizes of tokens) is 30. We also set the maximum number of tokens to 225, as the number of tokens can vary across images. Different from FH which relies on graph-based grouping, SLIC is similar to k-means, so an important hyper-parameter is the number of iterations used to optimize the segments.
The results for ImageNet classification are shown in Tab. 9. The patchification ($16\times 16$) performs best, even when compared against advanced tokens with additional training (e.g., FH does not converges as fast, so we used 400 epochs, longer than the default 300 epochs). SLIC (iter=1) performs close to patchification but visual inspection shows that it essentially reverts to patch-based tokens as part of the algorithm initialization. Further SLIC iterations, even with one more (iter=2), degrade performance. This suggests that the simple patchification is superior to more advanced grouping tokenization.
type of tokens Acc@1 Acc@5
patchification ($16\times 16$) 82.7 96.3
FH 81.3 95.6
SLIC (iter=2) 78.9 94.3
SLIC (iter=1) 82.4 96.1
Table 9: An exploration of advanced tokenizations. Here, we find that the performance improves when moving away from irregular shape and closer to patchification.